OpenStudy (anonymous):

simplify the equation by combining like terms 6x-2x=4

OpenStudy (anonymous):

bdboy0, all your questions today are solved the same way: Combine the x terms, then divide both sides by the coefficient of x.

OpenStudy (anonymous):

4x = 4 4x/4 = 4/4 x = 1

OpenStudy (anonymous):

im confused how did you get that???

OpenStudy (anonymous):

6 - 2 = 4. That means you have 4x's. 4x = 4 Divide each side by 4.. 4x/4 = 4/4 4/4 = 1 x = 1